OpenStudy (anonymous):

A high diver of mass 50.0 kg steps off a board 10.0 m above the water and falls vertical to the water, starting from rest. If her downward motion is stopped 2.20 s after her feet first touch the water, what average upward force did the water exert on her?

OpenStudy (shamim):

Let the velocity at first touch the water=v

OpenStudy (shamim):

Now required equation is v^2=u^2+2gh v^2=0^2+2*9.8*10 v=?

OpenStudy (shamim):

Negative acceleration within water is =a

OpenStudy (shamim):

Now we wanna use v=u-at 0=u-a*2.2 Here the value of u will b equal to the value of v calculated b4 a=?

OpenStudy (shamim):

Now use F=ma=?

OpenStudy (shamim):

Response plz for better understanding!!!"